OpenStudy (anonymous):

What is the standard form of the equation of the given circle with center A(4, 8)? A)(x - 4)^2 + (y - 8)^2 = 2.5 B)(x + 4)^2 + (y + 8)^2 = 6.25 C)(x + 4)^2 + (y + 8)^2 = 2.5 D)(x - 4)^2 + (y - 8)^2 = 6.25 PLZ EXPLAIN THX

OpenStudy (anonymous):



OpenStudy (anonymous):

@Abhisar @mathmale @math&ing001 please help

OpenStudy (anonymous):

i was thinkin it was A) but im not sure

OpenStudy (anonymous):

me to but this is tricky.

OpenStudy (anonymous):

but would guess A. tho

OpenStudy (anonymous):

ok thx 4 your help

OpenStudy (anonymous):

np

OpenStudy (anonymous):

:) good luck!

OpenStudy (math&ing001):

For a center A(4,8) and a radius of 2.5 the equation would be : (x - 4)^2 + (y - 8)^2 = (2.5)^2 => (x - 4)^2 + (y - 8)^2 = 6.25